OpenStudy (anonymous):

What formula do I use to determine the average rate of change for the function f(x)=3/4x-1?

OpenStudy (caozeyuan):

the derivative, f'=dy/dx=n*x^(n-1)

OpenStudy (lfreeman1285):

the -1 at the end would go away because the derivative of a constant is equal to zero.

OpenStudy (anonymous):

@caozeyuan is right. Use the derivative formula: \[ f'=\frac{ dy }{ dx }=n\ x ^{\left( n - 1 \right)}\] So it would look like this: \[f(x)=\frac{ 3 }{ 4 } x-1\] \[f'(x) = \frac{ 3 }{ 4 }\] The average rate of change of f(x) is 3/4.